When Samuel rides his bike at a speed of 22mph, it takes him 30 minutes to get home. Today he needs to get home in 25 minutes. How fast must he ride his bike to get home in time?
I get real flustered with these word problems. Is there an easy way to make it simple to understand?
Thanks!

Answer by stanbon(75887)

When Samuel rides his bike at a speed of 22mph, it takes him 30 minutes to get home. Today he needs to get home in 25 minutes. How fast must he ride his bike to get home in time?
-------------------------
Old DATA:
rate = 22 mph ; time = 1/2 hr ; distance = rt = (1/2)22 = 11 miles
--------------------------------------
New DATA:
rate = x mph ; time = (25/60) hr. distance = rt = (25/60)x miles
--------------------------------------
Equation:
new distance = old distance
(25/60)x = 11
x = (60/25)11 = 26.4 mph (speed required to go 11 miles in 25 minutes)
==========================================================================
Cheers,
Stan H.